OpenStudy (swaugh):

MEDAL + FANNN AWAITINGGGGG In a direct variation, y is -18 when x is 9. What is y when x equals -6? -12 -3 3 12

OpenStudy (michele_laino):

hint: direct variation between two variables \(x,y\), can be modeled by this formula: \(y=kx\)

OpenStudy (swaugh):

im confused

OpenStudy (michele_laino):

If I replace \(y=-18\) and \(x=9\), I get: \(-18=2\cdot k\) what is the value of \(k\)?

OpenStudy (swaugh):

-9

OpenStudy (michele_laino):

perfect! so, we can write this \(y=-2x\) now, please replace \(x=-6\), what is the value for \(y\)?

OpenStudy (michele_laino):

hint: \(y=-2x=-2 \cdot (-6)=...?\)

OpenStudy (michele_laino):

sorry I made an error, here is the right formula: \(-18=9 \cdot k\), so: \(k=-2\)

OpenStudy (swaugh):

i dont understand

OpenStudy (michele_laino):

hence: \[\huge y = - 2x = - 2 \cdot \left( { - 6} \right) = ...?\]

OpenStudy (michele_laino):

after I replace y=-18 and x=9, I get: \(-18=9k\) from which, the value of k is \(k=-2\) so, the formula fo direct variation is: \(y=-2x\) now, if I replace \(x=-6\) I get: \[y = - 2x = - 2 \cdot \left( { - 6} \right) = ...?\] so, what is the value of \(y\) ?

OpenStudy (michele_laino):

for*

OpenStudy (swaugh):

12

OpenStudy (swaugh):

wait is that the answer

OpenStudy (michele_laino):

that's right!

OpenStudy (swaugh):

thank you

OpenStudy (michele_laino):

it is the value you have found: \(y=12\)

OpenStudy (michele_laino):

:)